OpenStudy (anonymous):

Please help one question that i do not understand: Eliminate the parameter: x = 5 cos t, y = 5 sin t Please include steps

OpenStudy (anonymous):

I got this answer however i believe the final answer has to include an x. Please help? x ^2 + y^2 = (5cos t )^2 + (5sin t)^2 = 5^2 cos^2 + 5^2 sin^2 t = 5^2 ( cos^2 t + sin^2 t ) = 5^2 * 1

OpenStudy (anonymous):

I believe the final answer has to be y = something with an x on this side as well